# Performance and radiation damage mitigation strategy for silicon photomultipliers on LEO space missions

### v4

@@ Abstract
Space missions require lightweight, low-power consuming, radiation-tolerant components. Silicon photomultipliers are increasingly used for detecting near-UV, optical, and infrared light in space due to their compact design, low cost, low power consumption, robustness, and high photo-detection efficiency, which makes them sensitive to single photons. Although SiPMs outperform traditional photomultiplier tubes in many areas, concerns about their radiation tolerance and noise remain. In this study, we estimate the radiation effects on a satellite in sun-synchronous low Earth orbit (LEO) at an altitude of 550 km during the declining phase of solar cycle 25 (2026–2029). We evaluated silicon photomultipliers produced by the Foundation Bruno Kessler (FBK) using front-side illuminated technology with metal trenches (NUV-HD-MT), assessing their response to a 50 MeV proton beam and exposure to a -radioactive source (strontium-90). Simulations with SPENVIS and Geant4 were used to validate the experimental results. Based on our findings, we propose a photosensor annealing strategy for space-based instruments.
@@ 1 Introduction: silicon photosensors in space applications
The NUSES [5] project will employ SiPMs for both of its payloads, Terzina [6] and Ziré [7]. Ziré, dedicated to monitoring of low-energy ( MeV) cosmic ray fluxes in the Van Allen belts, gamma-ray bursts in the energy region 1-10 MeV, space weather and lithosphere-ionosphere- magnetosphere couplings, will employ SiPMs with LYSO or GAGG crystals in the calorimeter (CALOg), in the plastic scintillator tower (PST) reading scintillating tiles, and in the FTK fiber tracker.
@@ 2 The Terzina telescope onboard NUSES
The NUSES ballistic mission aims to explore new scientific and technological pathways for future space-based astroparticle physics detectors. The satellite platform, developed by the industrial partner TAS-I (THALES Alenia Space - Italia), will operate in a quasi-polar Low Earth Orbit (LEO) at an altitude of 535 in the worst case and 550 km at the Beginning of Life (BoL). The satellite will host two instruments: Terzina and Ziré. It will follow a sun-synchronous orbit and operate in a dusk-dawn mode along the day/night boundary so that the Terzina telescope can point to the dark side of the Earth’s limb to detect the faint Cherenkov light from ultra-relativistic particle showers in the dark atmosphere. The Terzina telescope aims at the first detection of this Cherenkov light, which can be detected within the field of view (FoV) of its camera located in the FPA. The FPA is designed to detect the Cherenkov light emitted by UHECR-induced EAS with energies above PeV and with their axis pointing upwards close to the optical axis of the telescope pointing to the limb.
@@
The Earth’s limb is at the angle , where when km at BoL. while at the end of life, when the drag of the residual atmosphere will reduce the orbit altitude to 535 km. In addition, Terzina can detect the upward EAS produced in the atmosphere by the decay of tau and muon leptons generated by the respective neutrinos with energies above PeV, passing through the Earth. A pictorial view of the detection principle of the Terzina payload on the NUSES satellite is presented in figure 1, highlighting the instrument’s positions and pointing directions, though not in scale. Terzina is a pathfinder for future larger missions, such as POEMMA or a constellations of satellites as conceived for CHANT [23].
@@
The Terzina detector consists of a near-UV optical telescope with Schmidt-Cassegrain optics (shown in figure 3 with effective collective area of 0.09 m2 and equivalent focal length of 925 mm and a FPA in the focal plane. The shadowing due to the support of the system (baffles and vanes) is 8%. The Optical Telescope Assembly (OTA), which will be realised by the company Officina Stellare, is composed of a hyperbolic primary mirror of 430 mm diameter and radius of curvature (RoC) of 1200 mm, fitting a flat FPA of 13160 mm2. The secondary aspherical mirror has diameter of 194 mm and RoC = 605 mm and it is located at a distance of 430 mm from the primary mirror. The secondary mirror is covered by a correcting lens of thickness 30 mm and RoC = 320 mm to compensate for all aberrations and principally for those induced by the flat FPA (if curved, it would have reduced aberrations but the mechanical design of the FPA would have been more complex).
@@
The FPA is designed to detect photons from both below and above the limb and consists of 10 SiPM arrays, each with pixels of mm2 (640 pixels in total). Each channel or pixel has a sensitive area of mm2. As it is known that a refractor and a Cassegrain optical system produce an inverted image when used without a diagonal mirror [24], the optical system of Terzina will image the Earth below the limb in the upper row of 5 SiPM arrays of the camera. These will be sensitive to eventual neutrino events from below the Earth’s limb, though the expected rate is negligble in Terzina. A larger mission will be required. They will be used for characterization of luminous noise from the Earth. The lower row of 5 SiPM arrays will observe EAS generated by UHECRs from above the limb. The telescope’s FoV is horizontally and vertically, as each pixel sees . A computer assisted design (CAD) model of the Terzina telescope is shown in figure 2 (on the left) and details of the CAD of the FPA and the tiles are shown on the right.
@@ 3 The NUV-HD-MT SiPMs for Terzina
In partnership with FBK, we have selected their technology minimizing the optical cross-talk and the DCR, the NUV-HD-MT [25]. We then investigated the micro-cell size that maximize the PDE while keeping the signal duration short. In fact, the signal duration is proportional to the micro-cell size and its capacitance. A large signal increases the noise captured within the acquisition window required to integrate the full pulse. The NUV-HD-MT technology has been developed by FBK adding metal-filled Deep Trench Isolation (DTI) to strongly suppress optical cross-talk. FBK finds a reduction in optical cross-talk by a factor of 10 [26] with respect to the NUV-HD in absence of the metall-filled DTI. We have measured the NUV-HD-MT SiPM of mm2 and mm2 with square micro-cell sizes of 25, 30, 35, 40, 50 . A further significant consideration is whether to utilise bare or coated sensors. While bare sensors require meticulous handling, the electrons may radiate Cherenkov light within the covering resin layer (see section 8).
@@ 3.1 Static characterisation
All the laboratory measurements have been performed at the University of Geneva, according to the setup and analysis methodology described in [4, 27]. Figure 4 shows the schematic layouts of the experimental setups for static (left) and dynamic/optical characterization (right).
The two configurations incorporate the following elements:
@@
For the static characterization configuration, the IV curve was measured by directly biasing and reading the SiPM using a Keithley 6487 or Keithley 2400 [29] picoammeter/voltage source. In the dynamic characterization configuration, the bias voltage was supplied by a Keithley 6487 picoammeter/voltage source reading before the signal with a 2 GHz bandwidth oscilloscope. A custom-made amplifier with a bandwidth of approximately MHz and a gain of dB was used to amplify, pulse shape and filter the SiPM signal, and effectively suppressing the slow component of the SiPM response. The Keithley 2000 multimeter [30], shown in this configuration, has been used for reading a temperature sensor mounted near the control photodiode. The average temperature during the measurements was ∘C (hereinafter referred to as room temperature).
@@ 3.1.1 Forward IV characterisation
The forward IV characteristic curve of the FBK SiPMs, shown in figure 5 (left panel), exhibits a very small increase in current when the polarization voltage, , is below the threshold value, followed by a rapid, linear increase in current with above this threshold. This behaviour can be described by the Shockley law, which governs the forward current , flowing through a junction diode. A SiPM is an array of -cells () connected in parallel, each referred to as a SPAD (single photon avalanche diode). Each -cell can be modelled as a diode in series with a quenching resistor, . For the entire device, the voltage can be expressed as:
@@ 3.2 Dynamic characterisation
To perform the dynamic characterisation of SiPM and measure the uncorrelated noise, we utilized the setup shown in the right panel of Figure 4. A full over-voltage scan was conducted by ranging the voltage between 1 V and 18 V, and by acquiring 10’000 waveforms on the oscilloscope for each given over-voltage. To compute the DCR and OCT, the analysis specifically focused on the signals received within dark interval of the waveform (in the absence of light).
@@ 3.3 Optical characterisation
The photon detection efficiency (PDE) is a crucial property of SiPMs, quantifyng how effectively the device detects incoming photons as a function of their wavelength (). The PDE can be described as:
@@
Where is the quantum efficiency, i.e. the probability that a photon generates an electron-hole pair, is the fill factor, representing the percentage of the sensor area sensitive to light333The fill factor is influenced by the -cell sizes and their effective area at the pixel level. Specifically, the fill factor depends on how much of the total pixel area is occupied by the -cells compared to non-active regions (such as gaps or interconnects) within the pixel. and is the Geiger probability. The PDE strongly depends on the wavelength of the incident light, the over-voltage, and consequently, the breakdown voltage. In addition, PDE is temperature-dependent, as carrier mobility increases with temperature, potentially leading to an increase in PG for a fixed over-voltage. However, at around room temperature, this effect is likely negligible. Nonetheless, if the over-voltage is not adjusted to compensate for temperature variations, PDE measurements will fluctuate along with the breakdown voltage.
@@
Using the same setup as in the dynamic characterization (right panel of Figure 4), we performed a calibration step by replacing the SiPM with a calibrated control photodiode to determine the absolute amount of incoming light. After reinstalling the FBK SiPM, we repeated the scan while illuminating it with a low-intensity pulsed LED. Following the analysis methodology of [27], we focused on the trigger window (the expected arrival time of photons from the LED). The PDE was calculated using the Poisson method, based on the average number of detected photons and corrected for uncorrelated SiPM noise. Figure LABEL:fig:pde_all_b shows the PDE as a function of the cell size for different over-voltages and sensor coatings. By looking at this plot and considering the results of the OCT and DCR, the 30 m cell size bare sensor appears to be the best solution for a LEO satellite. At 10 V of over-voltage, in fact, the PDE for 30 m cell size is already more than 50 (see figure LABEL:fig:pde_all_a).
@@
Weighing the DCR and PDE, as described above, against the pulse shape as outlined in [8], it has been decided to use SiPM tiles with a 30 m bare -cell size. At this configuration, we have an optimal performance in DCR and PDE, whilst still having a short pulse decay time. The decay time and the PDE are anti-correlated: smaller -cell sizes result in shorter signal durations but also in lower PDE due to the reduced fill factor. Hence, the final choice of the SiPM -cell of 30 m represents a trade-off between these parameters. Additionally it complies with the constraints on the maximum power consumption of the NUSES satellite. As a matter of fact, the larger the DCR, the larger the power consumption of the SiPM power supplies. If the data acquisition thresholds are not adjusted accordingly, the rate of noise induced triggers would also require more processing activity, further increasing power consumption. We performed a set of measurements with three different configurations: 1) with a pulsed LED (450 nm) and an integration sphere; 2) with a short-duration (25 ps) light-pulse laser with a 370 nm wavelength laser with and 3) without the integration sphere. Both the LED and the laser flash the full sensor to obtain a high amplitude signal to be detectable by a 2 GHz bandwidth oscilloscope. Figure 11 summarizes the obtained results. As expected, the shortest decay time was measured with the laser only, as the integration sphere induces an additional time spread. The SiPM signal tail has been fitted with a negative exponential function and the decay time is in the figure. This time is a function of the -cell capacitance, hence its size. We did not observe any significant change in the signal shape with the variation in the bias voltage.
@@ 4 Evaluation of radiation fluxes for a LEO space mission
To estimate the total radiation we considered solar particle events (SPEs), trapped protons and electrons in the Van Allen radiation belts and galactic cosmic rays (GCRs). These expected fluxes for the NUSES spacecraft averaged on 15 orbits (1 day) during its operation time are shown in figure 13. The shaded areas indicate the fluctuations of these fluxes depending on the variation of the satellite altitude between 500 and 600 km and considering a period of maximum solar activity. In particular, the higher the altitude, the higher the flux of trapped protons and trapped electrons and consequently the slightly higher the flux of solar protons. Since the flight will occur during the declining part of the solar maximum, the fluxes for the trapped protons and electrons have been calculated using the NASA models AP-8 max and AE-8 max [35, 36], respectively, and the solar protons using the magnetic shield on in SAPPHIRE [37], and ISO 15390 model [38] for the GCRs. We used the realistic case of 550 km. It can be noticed that, for our altitude, the GCR contribution (protons and helium nuclei)444Heavier nuclei than p GCR fluxes are much lower at these energies and are neglected in this analysis. is negligible. Moreover, since NUSES will operate during intense solar activity, the solar flux will be very high while the galactic proton flux will be reduced [39] as the solar wind impedes GCR penetration into the solar system. The main contributions to the radiation damage to the camera will be due to trapped protons and electrons, and solar protons. Although it is partially accounted for in the trapped particle models, Earth’s albedo has not been evaluated separately because can be considered negligible for a LEO mission.
@@ 5 Dose estimates from radiation in space with Geant4
Protons and electrons have been injected into these geometries, simulated in Geant4, to estimate the total radiation dose. Particles were simulated in small energy bins considering the full energy range of the flux calculated with SPENVIS integrated over a sphere of radius cm surrounding the considered volumes 555It is important to note that, in this study, the flows were considered to be isotropic, which is a conservative case compared to the real scenario. The direction and regions where we expect the presence of these particles have been produced in SPENVIS.and for the duration of the mission of 3 years. The total energy deposited in the sensitive volume of the FPA tiles and in the PCB, returned in MeV, has been calculated by Geant4 666Secondary particle production has been simulated and taken into account. Material activation, on the other hand, was considered to be a minor concern and negligible in terms of the total dose.. The total dose released in the corresponding volumes for all particles in the energy bin has been calculated by summing all the differential doses:

### v5

@@ Abstract
Space missions require lightweight, low-power consuming, radiation-tolerant components. Silicon photomultipliers (SiPMs) are increasingly used for detecting near-UV, optical, and infrared light in space due to their compact design, low cost, low power consumption, robustness, and high photo-detection efficiency, which makes them sensitive to single photons. Although SiPMs outperform traditional photomultiplier tubes in many areas, concerns about their radiation tolerance and noise remain. In this study, we estimate the radiation effects on a satellite in sun-synchronous low Earth orbit (LEO) at an altitude of 550 km during the declining phase of solar cycle 25 (2026–2029). We evaluated silicon photomultipliers produced by the Foundation Bruno Kessler (FBK) using front-side illuminated technology with metal trenches (NUV-HD-MT), assessing their response to a 50 MeV proton beam and exposure to a -radioactive source (strontium-90). Simulations with SPENVIS and Geant4 were used to validate the experimental results. Based on our findings, we propose a photosensor annealing strategy for space-based instruments.
@@ 1 Introduction: silicon photosensors in space applications
The NUSES [5] project will employ SiPMs for both of its payloads, Terzina [6] and Ziré [7]. Ziré, dedicated to monitoring of low-energy ( MeV) cosmic ray fluxes in the Van Allen belts, gamma-ray bursts in the energy region 1-10 MeV, space weather and lithosphere-ionosphere- magnetosphere couplings, will employ SiPMs with LYSO or GAGG crystals in the calorimeter (CALOg), in the plastic scintillator tower (PST) reading scintillating tiles, and in the fiber tracker (FTK).
@@ 2 The Terzina telescope onboard NUSES
The NUSES ballistic mission aims to explore new scientific and technological pathways for future space-based astroparticle physics detectors. The satellite platform, developed by the industrial partner TAS-I (THALES Alenia Space - Italia), will operate in a quasi-polar Low Earth Orbit (LEO) at an altitude of 535 km in the worst case and 550 km at the Beginning of Life (BoL). The satellite will host two instruments: Terzina and Ziré. It will follow a sun-synchronous orbit and operate in a dusk-dawn mode along the day/night boundary so that the Terzina telescope can point to the dark side of the Earth’s limb to detect the faint Cherenkov light from ultra-relativistic particle showers in the dark atmosphere. The Terzina telescope aims at the first detection of this Cherenkov light, which can be detected within the field of view (FoV) of its camera located in the FPA. The FPA is designed to detect the Cherenkov light emitted by UHECR-induced EAS with energies above PeV and with their axis pointing upwards close to the optical axis of the telescope pointing to the limb.
@@
The Earth’s limb is at the angle , where when km at BoL while at the end of life, when the drag of the residual atmosphere will reduce the orbit altitude to 535 km. In addition, Terzina can detect the upward EAS produced in the atmosphere by the decay of tau and muon leptons, which are generated by the respective neutrinos with energies above PeV, passing through the Earth. A pictorial view of the detection principle of the Terzina payload on the NUSES satellite is presented in figure 1, highlighting the instrument’s positions and pointing directions, though not in scale. Terzina is a pathfinder for future larger missions, such as POEMMA or a constellation of satellites as conceived for CHANT [23].
@@
The Terzina detector consists of a near-UV optical telescope with Schmidt-Cassegrain optics (shown in figure 3) with effective collective area of 0.09 m2 and equivalent focal length of 925 mm and a FPA in the focal plane. The shadowing due to the support of the system (baffles and vanes) is 10%. The Optical Telescope Assembly (OTA), which will be realised by the company Officina Stellare, is composed of a hyperbolic primary mirror of 430 mm diameter and radius of curvature (RoC) of 1200 mm, fitting a flat FPA of 13160 mm2. The secondary aspherical mirror has a diameter of 194 mm and RoC = 605 mm and it is located at a distance of 430 mm from the primary mirror. The secondary mirror is covered by a correcting lens of thickness 30 mm and RoC = 320 mm to compensate for all aberrations and principally for those induced by the flat FPA (if curved, it would have reduced aberrations but the mechanical design of the FPA would have been more complex).
@@
The FPA is designed to detect photons from both below and above the limb and consists of 10 SiPM arrays, each with pixels of mm2 (640 pixels in total). Each pixel has a sensitive area of mm2. As it is known that a refractor and a Cassegrain optical system produce an inverted image when used without a diagonal mirror [24], the optical system of Terzina will image the Earth below the limb in the upper row of 5 SiPM arrays of the camera. These will be sensitive to potential neutrino events from below the Earth’s limb, though the expected rate is negligible for Terzina. A larger mission will be required. They will be used for characterisation of luminous noise from the Earth. The lower row of 5 SiPM arrays will observe EAS generated by UHECRs from above the limb. The telescope’s FoV is horizontally and vertically, as each pixel sees . A computer assisted design (CAD) model of the Terzina telescope is shown in figure 2 (on the left) and details of the CAD of the FPA and the tiles are shown on the right.
@@ 3 The NUV-HD-MT SiPMs for Terzina
In partnership with FBK, we have selected their technology minimizing the optical cross-talk and the DCR, the NUV-HD-MT [25]. We then investigated which micro-cell size would maximise the PDE while keeping the signal duration short. In fact, the signal duration is proportional to the micro-cell size and its capacitance. A large signal increases the noise captured within the acquisition window required to integrate the full pulse. The NUV-HD-MT technology has been developed by FBK adding metal-filled Deep Trench Isolation (DTI) to strongly suppress optical cross-talk. FBK finds a reduction in optical cross-talk by a factor of 10 [26] with respect to the NUV-HD in absence of the metal-filled DTI. We have measured the NUV-HD-MT SiPM of mm2 and mm2 with square micro-cell sizes of 25, 30, 35, 40, 50 . Another significant consideration is whether to use bare or coated sensors. While bare sensors require meticulous handling, the electrons may radiate Cherenkov light within the covering resin layer (see section 8).
@@ 3.1 Static characterisation
All the laboratory measurements have been performed at the University of Geneva, according to the setup and analysis methodology described in [4, 27]. Figure 4 shows the schematic layouts of the experimental setups for static (left) and dynamic/optical characterisation (right). The two configurations incorporate the following elements:
@@
For the static characterisation configuration, the IV curve was measured by directly biasing and reading the SiPM using a Keithley 6487 or Keithley 2400 [29] picoammeter/voltage source. In the dynamic characterisation configuration, the bias voltage was supplied by a Keithley 6487 picoammeter/voltage source reading before the signal with a 2 GHz bandwidth oscilloscope. A custom-made amplifier with a bandwidth of approximately MHz and a gain of dB was used to amplify, pulse shape and filter the SiPM signal, and effectively suppress the slow component of the SiPM response effectively. The Keithley 2000 multimeter [30], shown in this configuration, has been used for reading a temperature sensor mounted near the control photodiode. The average temperature during the measurements was ∘C (hereinafter referred to as room temperature).
@@ 3.1.1 Forward IV characterisation
The forward IV characteristic curve of the FBK SiPMs, shown in figure 5 (left panel), exhibits a very small increase in current when the polarization voltage, , is below the threshold value, followed by a rapid, linear increase in current with above this threshold. This behaviour can be described by the Shockley law, which governs the forward current , flowing through a - junction diode. A SiPM is an array of -cells () connected in parallel, each referred to as a SPAD (single photon avalanche diode). Each -cell can be modelled as a diode in series with a quenching resistor, . For the entire device, the voltage can be expressed as:
@@ 3.2 Dynamic characterisation
To perform the dynamic characterisation of SiPM and measure the uncorrelated noise, we used the setup shown in the right panel of figure 4. A full over-voltage scan was conducted by ranging the voltage between 1 V and 18 V, and by acquiring 10’000 waveforms on the oscilloscope for each given over-voltage. To compute the DCR and OCT, the analysis specifically focused on the signals received within dark interval of the waveform (in the absence of light).
@@ 3.3 Optical characterisation
The photon detection efficiency (PDE) is a crucial property of SiPMs, quantifying how effectively the device detects incoming photons as a function of their wavelength (). The PDE can be described as:
@@
Where is the quantum efficiency, i.e. the probability that a photon generates an electron-hole pair, is the fill factor, representing the percentage of the sensor area sensitive to light333The fill factor is influenced by the -cell sizes and their effective area at the pixel level. Specifically, the fill factor depends on how much of the total pixel area is occupied by the -cells compared to non-active regions (such as gaps or interconnects) within the pixel. and is the Geiger probability. The PDE strongly depends on the wavelength of the incident light, the over-voltage, and consequently, the breakdown voltage. In addition, PDE is temperature-dependent, as carrier mobility increases with temperature, potentially leading to an increase in for a fixed over-voltage. However, at around room temperature, this effect is likely negligible. Nonetheless, if the over-voltage is not adjusted to compensate for temperature variations, PDE measurements will fluctuate along with the breakdown voltage.
@@
Using the same setup as in the dynamic characterisation (right panel of figure 4), we performed a calibration step by replacing the SiPM with a calibrated control photodiode to determine the absolute amount of incoming light. After reinstalling the FBK SiPM, we repeated the scan while illuminating it with a low-intensity pulsed LED. Following the analysis methodology of [27], we focused on the trigger window (the expected arrival time of photons from the LED). The PDE was calculated using the Poisson method, based on the average number of detected photons and corrected for uncorrelated SiPM noise. Figure LABEL:fig:pde_all_b shows the PDE as a function of the cell size for different over-voltages and sensor coatings. By looking at this plot and considering the results of the OCT and DCR, the 30 m cell size bare sensor appears to be the best solution for a LEO satellite. At 10 V of over-voltage, in fact, the PDE for 30 m cell size is already more than 50 (see figure LABEL:fig:pde_all_a).
@@
Weighing the DCR and PDE, as described above, against the pulse shape as outlined in [8], it has been decided to use SiPM tiles with a 30 m bare -cell size. With this configuration, we achieve optimal performance in terms of DCR and PDE while maintaining a short pulse decay time. The decay time and the PDE are anti-correlated: smaller -cell sizes result in shorter signal durations but also in lower PDE due to the reduced fill factor. Hence, the final choice of the SiPM -cell of 30 m represents a trade-off between these parameters. Additionally it complies with the constraints on the maximum power consumption of the NUSES satellite. As a matter of fact, the larger the DCR, the larger the power consumption of the SiPM power supplies. If the data acquisition thresholds are not adjusted accordingly, the rate of noise induced triggers would also require more processing activity, further increasing power consumption. We performed a set of measurements with three different configurations: 1) with a pulsed LED (450 nm) and an integration sphere; 2) with a short-duration (25 ps) light-pulse laser with a 370 nm wavelength laser with and 3) without the integration sphere. Both the LED and the laser flash the full sensor to obtain a high amplitude signal to be detectable by a 2 GHz bandwidth oscilloscope. Figure 11 summarises the obtained results. As expected, the shortest decay time was measured with the laser only, as the integration sphere induces an additional time spread. The SiPM signal tail has been fitted with a negative exponential function and the decay time is in the figure. This time is a function of the -cell capacitance, hence its size. We did not observe any significant change in the signal shape with the variation in the bias voltage.
@@ 4 Evaluation of radiation fluxes for a LEO space mission
To estimate the total radiation we considered solar particle events (SPEs), trapped protons and electrons in the Van Allen radiation belts and galactic cosmic rays (GCRs). These expected fluxes for the NUSES spacecraft averaged on 15 orbits (1 day) during its operation time are shown in figure 13. The shaded areas indicate the fluctuations of these fluxes depending on the variation of the satellite altitude between 500 km and 600 km and considering a period of maximum solar activity. In particular, the higher the altitude, the higher the flux of trapped protons and trapped electrons and consequently the slightly higher the flux of solar protons. Since the flight will occur during the declining part of the solar maximum, the fluxes for the trapped protons and electrons have been calculated using the NASA models AP-8 max and AE-8 max [34, 35], respectively, and the solar protons using the magnetic shield on in SAPPHIRE [36], and ISO 15390 model [37] for the GCRs. We used the realistic case of 550 km. It can be noticed that, for our altitude, the GCR contribution (protons and helium nuclei)444Heavier nuclei than p GCR fluxes are much lower at these energies and are neglected in this analysis. is negligible. Moreover, since NUSES will operate during intense solar activity, the solar flux will be very high while the galactic proton flux will be reduced [38] as the solar wind impedes GCR penetration into the solar system. The main contributions to the radiation damage to the camera will be due to trapped protons and electrons, and solar protons. Although it is partially accounted for in the trapped particle models, Earth’s albedo has not been evaluated separately because can be considered negligible for a LEO mission.
@@ 5 Dose estimates from radiation in space with Geant4
Protons and electrons have been injected into these geometries, simulated in Geant4, to estimate the total radiation dose. Particles were simulated in small energy bins considering the full energy range of the flux calculated with SPENVIS integrated over a sphere of radius cm surrounding the considered volumes 555It is important to note that, in this study, the fluxes were considered to be isotropic, which is a conservative case compared to the real scenario. The direction and regions where we expect the presence of these particles have been produced in SPENVIS.and for the duration of the mission of 3 years. The total energy deposited in the sensitive volume of the FPA tiles and in the PCB, returned in MeV, has been calculated by Geant4 666Secondary particle production has been simulated and taken into account. Material activation, on the other hand, was considered to be a minor concern and negligible in terms of the total dose.. The total dose released in the corresponding volumes for all particles in the energy bin has been calculated by summing all the differential doses: